OpenStudy (anonymous):

PLEASE HELP ME !!!!!!!!!!! What value of b makes the statement true? 17 = b – (–6.2) A. –23.2 B. –10.8 C. 10.8 D. 23.2

OpenStudy (shinalcantara):

17 = b - (-6.2) removing the parenthesis, the negative of a negative number is positive making it 17 = b + 6.2 subtract both sides of the equation by 6.2 17-6.2 = b+6.2-6.2 b=? evaluate

OpenStudy (anonymous):

i could help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Start by Simplifying 17 = b + -1(-6.2) Multiply -1 * -6.2 17 = b + 6.2 Then Reorder the terms: 17 = 6.2 + b Solving 17 = 6.2 + b Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add '-1b' to each side of the equation. 17 + -1b = 6.2 + b + -1b Next, Combine like terms: b + -1b = 0 17 + -1b = 6.2 + 0 17 + -1b = 6.2 Add '-17' to each side of the equation. 17 + -17 + -1b = 6.2 + -17 Combine like terms: 17 + -17 = 0 0 + -1b = 6.2 + -17 -1b = 6.2 + -17 Combine like terms: 6.2 + -17 = -10.8 -1b = -10.8 Divide each side by '-1'. b = 10.8 Simplify b = 10.8 Now your done! Hope that helped!

OpenStudy (shinalcantara):

what did you get?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

@curiousdreamer

OpenStudy (anonymous):

D is incorrect.

OpenStudy (anonymous):

really

OpenStudy (anonymous):

yes your right its b

OpenStudy (anonymous):

B is incorrect also

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

wat i suk

OpenStudy (anonymous):

thx sunnyshores

OpenStudy (anonymous):

No. You just did something wrong with your calculating. it happens all the time.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.